OpenStudy (anonymous):

Please help! Use a sketch to find the exact value of the expression: sin[cos^-1(square root 2/2)]

OpenStudy (apoorvk):

\[\frac {\sqrt 2}{2} = \frac {1}{\sqrt 2}\] what's cos^-1(1/sqrt2) ?

OpenStudy (anonymous):

using the calculator I get .7854, is this the exact value for the question?

OpenStudy (apoorvk):

do you know that sin(45degrees) = cos(45degrees) = 1/sqrt2 ?? This is one of the standard values you should remember actually.

OpenStudy (anonymous):

I do, but it has been many years and I am trying to remember the process.

OpenStudy (apoorvk):

Well, cos or sin inverse of x is that angle for which the cos or sin value is x.

OpenStudy (anonymous):

So then the angle on the unit circle corresponds to \[\pi/4\]

OpenStudy (apoorvk):

Yes, exactly!!!

OpenStudy (anonymous):

Great! Thanks!

OpenStudy (apoorvk):

No worries! :)